What is the total electric flux through a closed surface containing a 2.0 μC charge?



1
Expert's answer
2022-04-06T18:36:29-0400

The Gauss's law says

"\phi=\oint EdA=\frac{q}{\epsilon_0}"

"\phi=\frac{2.0*10^{-6}}{8.85*10^{-12}}=2.3*10^5\:\rm N\cdot m^2/C"